OpenStudy (anonymous):

What are the x and y intercepts of y = 2(x+3)^2 - 1?

OpenStudy (anonymous):

x-intercept is when y = 0 y-intercept is when x = 0

OpenStudy (utterly_confuzzled):

You can get the y-intercept by substituting 0 for x You can get the x-intercept by putting the equation equal to zero and then either factoring, using the quadratic formula, or guessing and checking. Since it's a quadratic equation, there will be 2 x-intercepts, so make sure if you choose to guess and check you find 2 numbers that work

OpenStudy (anonymous):

y intercept is (0, 17).

OpenStudy (anonymous):

I couldn't find the x intercept though? Y doesn't touch 0.

OpenStudy (anonymous):

Are these right?

OpenStudy (utterly_confuzzled):

0= 2(x+3)^2 - 1 0=2(x^2+6x+9)-1 0=2x^2+12x+17 \[\frac{ (-12)\pm \sqrt{12^{2}-4(2)(17} }{ 2(2) }\]

OpenStudy (utterly_confuzzled):

That should gve you the 2 x intercepts

OpenStudy (anonymous):

\[\frac{ -12\pm \sqrt{8} }{ 4 }\] (-2.3, 0), (-3.7, 0)?

OpenStudy (utterly_confuzzled):

You could also just keep them in the form with the square root (just be sure to divide by 4 so that it's in its simplest form).

OpenStudy (anonymous):

Okay, but the answers i put would be right?

OpenStudy (utterly_confuzzled):

Yup

OpenStudy (anonymous):

Thanks, if you don't mind, would you help me with the intervals increasing and decreasing?

OpenStudy (utterly_confuzzled):

Sure! It may be a bit slow though, I'm cooking....

OpenStudy (utterly_confuzzled):

The vertex is -3, so your decreasing interval is negative infinity ----> -3 And you increasing interval is -3 -------> infinity

OpenStudy (anonymous):

Thank you!